Suppose the end of your finger is charged. Estimat
Answer #1

Solution:

a) Air starts conducting when the electric field exceeds 3 x 10^6 N/C.

Finger tip is assumed to be a part of a sphere of diameter 1 cm.,(region surrounding close to the sphere is assumed to be spherical)..

Break down Voltage = V = E * R = (3 x 10^6) ( 0.005) = 15000 Volts.

Since this is only an estimate it can be rounded to 20,000 volts = 2 x 10^4 V( 1 significant figure)

b) Surface charge density = Q/A = charge /area

Charge = (4 pi eo R) V since V = 1/ 4 pi eo * Q/R

So Surface charge density = \sigma = ( 4 pi eo R) V / 4 pi R^2 = eo V/R = (8.85 x10^-12)(20000) / (0.005)

= 0.00000354 C/m^2

= 4 \mu C/m^2 (one significant figure)